OpenStudy (anonymous):

[9.02] What are the solutions of 2x2 = 24x? x = 0, x = 12 x = 0, x = −12 x = 2, x = 12 x = 2, x = −12

OpenStudy (anonymous):

First, I can tell you that the answer is x=0, x=12.

OpenStudy (anonymous):

hows that

OpenStudy (anonymous):

2x2=24x 2x2-24x=0 2(x2-12x)=0 x2-12x=0 x(x-12)=0 You see, if x=0, the whole sentence becomes 0; the other situation is that (x-12)=0, so x=12, and the whole sentence will be 0 too. I hope I explain it well.

OpenStudy (anonymous):

yes you did thank you so much!!!!!!!!!!

OpenStudy (anonymous):

You are welcome. I am glad to answer for you.

OpenStudy (anonymous):

:)